Maggie0803:

What are the solutions to the equation? x2−14x=15 Enter your answers in the boxes. x = ____ and ____

misty34:

Add 14 to 15 and what do you get?

jhonyy9:

misty this is not possibile bc. -14 is an x term

misty34:

Oops. Wrong thing, sorry

jhonyy9:

like a first step substract from both sides 15

misty34:

That's for regular equation

Maggie0803:

Add 14 to both sides right?

misty34:

It's been a while since i've done these

dude:

misty, he's right D:

misty34:

I know, I meant I was thinking of the way to do regular equations, sorry, I segmented it

jhonyy9:

Maggie do you understand ?

jhonyy9:

x^2 -14x -15 = 15-15

misty34:

Put it in standard form then use quadratic equation

Maggie0803:

Yeah, I think

jhonyy9:

no please make what i ve wrote above firstly

jhonyy9:

bc. after this step you will see that will can factorizing completly what will get

jhonyy9:

Maggie please collaborate

jhonyy9:

please continue x^2 -14x -15 = 15 -15

misty34:

Quadratic formula I meant

jhonyy9:

not is necessary use the quadratic formula here

misty34:

Use this formula? x = −b ± √(b2 − 4ac)

jhonyy9:

x^2 -14x -15 = 0 do you know two numbers with sum equal -14 and product equal -15 ?

Maggie0803:

It would be 15 and -1, right?

jhonyy9:

not exactly

jhonyy9:

15-1 not equal -14

Maggie0803:

15-1 is +14

jhonyy9:

yes but you need getting -14 assuming these two numbers

jhonyy9:

courage Maggie this is easy

misty34:

-14-1?

misty34:

-13

jhonyy9:

misty how is this right ? you need getting -14

misty34:

I meant -13-1

jhonyy9:

misty -14-1 = -15not -13

jhonyy9:

ok but the product need being -15

Maggie0803:

So -15 and -1?

jhonyy9:

how ? -15-1 = -16

misty34:

-13-1=-14

Maggie0803:

This is so confusing

jhonyy9:

misty ok but for product you need getting -15

misty34:

-14-1=-15

jhonyy9:

product -15 and sum -14

misty34:

x2−14x=15x2-14 x=15. Move 15 15 to the left side of the equation by subtracting it from both sides.

dude:

Alright, just to set some clarity here jhonny is guiding you through factoring, which is an alternative to the quadratic equation You want to find 2 numbers that multiply to -15 and add up to 14 The factors of 15 are (Factors: numbers that multiply to a specific one) 1,15 Now we know one has to be negative since 15 is negative (x-)(x+) Which numbers add up to -14? (one having to be negative) 1 and -15 Hence we should have (x+1)(x-15)=0

Maggie0803:

Thx u

dude:

You should set each factor equal to 0 to find your zeroes \(x\color{red}{\cancel{+1}}=0\) and \(x \color{red}{\cancel{-15}}=0\) \(\color{red}{\cancel{-1}}~~\color{red}{-1}\) \(\color{red}{\cancel{+15}}~~\color{red}{+15}\) Do the opposite of each number, (if it is adding, subtract; likewise if it is subtracting, add) \(x= -1\) and \(x=15\) Am late but was trying to get the latex through :c